# MemWeaver: A Hierarchical Memory from Textual Interactive Behaviors for Personalized Generation

### v1

@@ 4.4.2. Analysis of Incremental Update Strategy
To evaluate the practicality of our incremental update strategy in dynamic environments, we designed an experiment simulating the sequential arrival of user data. While maintaining up-to-date user profiles is critical, constantly reconstructing the entire graph is computationally prohibitive. As shown in Figure 5, our experiment compares the accuracy and efficiency of three distinct strategies: a full rebuild of the graph, a baseline with no updates, and our proposed incremental update approach. The two baseline strategies define the boundaries of the performance-efficiency spectrum. The full rebuild approach serves as an accuracy ceiling, achieving the highest possible performance by leveraging all available data for reconstruction. However, its high computational cost, depicted by the flat line at the bottom of the efficiency scale, renders it impractical for real-time applications. Conversely, the no-update strategy represents the efficiency ceiling by incurring no update cost. This efficiency comes at a steep price, however, as its accuracy degrades substantially when using a stale or incomplete user profile.
Our proposed incremental update strategy successfully navigates this trade-off. In terms of accuracy, it closely tracks the optimal performance of a full rebuild. The minor accuracy gap, which is only noticeable when updating from a very sparse initial history, diminishes as the initial graph becomes more complete, ultimately converging with the full rebuild performance. Simultaneously, its efficiency is substantially higher than that of a full rebuild and remains much closer to the near-instantaneous no-update baseline, demonstrating the remarkably low computational overhead of our approach. In conclusion, the experimental results validate that our incremental update mechanism provides a highly effective and practical solution. It resolves the critical trade-off between model accuracy and computational efficiency, delivering performance nearly indistinguishable from a complete graph reconstruction at a fraction of the cost. This capability is crucial for deploying sophisticated and responsive user models in dynamic, real-world systems.
@@ 4.6. Hyperparameter Sensitivity Analysis
To understand the mechanisms of MemWeaver, we analyze key hyperparameters in behavioral memory construction as shown in Figure 6, focusing on semantic guidance, and temporal biases, to reveal their impact on context retrieval.
@@ 5. Conclusion
In this study, we propose MemWeaver, a novel framework that addresses the challenge of deep personalization from unstructured user histories by restructuring them into a hierarchical memory. The framework’s core innovation is the effective interaction between its two memory components. The behavioral memory grounds the model in specific, contextual actions, and the cognitive memory provides abstract guidance on long-term preferences. This powerful combination provides a comprehensive representation for generating content deeply aligned with a user’s latent preferences. Crucially, the framework also supports efficient incremental updates, ensuring its scalability and responsiveness in dynamic applications. Extensive experiments on the six datasets of the LaMP benchmark demonstrate the efficacy of this approach. We hope our framework offers new inspiration for the future of user behavior modeling.
@@ Appendix B Dataset Details
Table 4. Prompt structure for generating a Local Summary from a behavior cluster.
Component Content / Description
System Role ”You are an expert at analyzing user behavior patterns.”
Input Context A numbered list of all user activities within a single cluster, prefaced with:
**Cluster X Activities (Y records):**
Task Instruction ”Provide a concise summary of the user’s preferences and behavior patterns in this cluster. Focus on key themes, preferences, and patterns. Use clear, structured language.”
Table 5. Prompt structure for synthesizing Local Summaries into a Global Summary (Cognitive Memory).
Component Content / Description
System Role ”You are an expert at creating concise user preference summaries.”
Input Context A list of all previously generated Local Summaries, formatted as:
Cluster 1: [Text of local summary 1]
Cluster 2: [Text of local summary 2]
Task Instruction ”Create a concise global summary (max 300 words) that captures the user’s key preferences and behavior patterns. Focus on the most important themes and avoid redundancy. Use bullet points or short paragraphs for clarity. Be brief but comprehensive.”
@@
LaMP-7: Personalized Tweet Paraphrasing
This task requires paraphrasing a given tweet in a style consistent with the user’s historical tweets. The data is from the Sentiment140 dataset. The input is a tweet that needs to be paraphrased. The user’s profile is a list of their previous tweets. The output is the paraphrased tweet, reflecting the user’s unique style.
@@
Appendix C Prompt Templates
This section details the prompt structures used within the MemWeaver framework. We first present the prompts used in the two-stage process for constructing the cognitive memory. Following that, we describe the unified template for Memory-Augmented Generation, which utilizes the constructed memories to perform personalized tasks on the LaMP benchmark. The design of these prompt templates aims to ensure consistency, controllability, and interpretability of the LLM’s reasoning process. By explicitly structuring role definitions, input contexts, and task instructions, MemWeaver standardizes how the model extracts user-level knowledge and applies it to downstream personalized tasks. Each component is carefully optimized to balance informativeness and prompt length, avoiding over-conditioning or unnecessary verbosity that may harm generalization across different task types.
C.1. Prompt for Cognitive Memory
The cognitive memory abstracts long-term user preferences and global behavior trends through a two-stage summarization process. First, the model produces a Local Summary for each cluster of user activities grouped by temporal or semantic similarity, capturing local coherence while filtering redundant interactions. Then, these local summaries are aggregated into a concise Global Summary that encapsulates the user’s stable cognitive profile and higher-level preference tendencies. This hierarchical process mirrors human cognition—observing specific behaviors before abstracting enduring traits. Both stages are guided by carefully designed prompts (Table 4, Table 5) that encourage concise, interpretable summaries emphasizing semantic generalization over repetition.
C.2. Prompt for Personalized Generation
After constructing the cognitive and behavioral memories, our MemWeaver uses a unified personalized generation prompt that integrates both sources to guide reasoning and output generation. The cognitive memory provides high-level priors reflecting long-term preferences, while the behavioral memory offers concrete cues relevant to the current query. Together, they act as complementary signals—global context and local evidence—that steer the model toward user-aligned outputs. As shown in Table 6 (LaMP-1 example), this modular prompt structure prepends the memory components before the query, ensuring that personalization cues are explicitly incorporated during generation. Its consistent design improves interpretability and transferability across diverse tasks.
Appendix D Additional Experimental Results
In this section, we conduct further experiments to provide a deeper analysis of our MemWeaver framework. We specifically investigate two key aspects: the impact of using different underlying embedding models, and the performance scalability as the size of the backbone language model increases.
D.1. Impact of Different Embedding Models
@@
To verify that the performance gains of our framework are not solely dependent on a specific embedding model, we conducted a comprehensive experiment to evaluate its robustness across different retrieval backbones. We replaced the default retriever in both the standard RAG baseline and our MemWeaver framework with three distinct models: a sparse retriever (BM25), a strong dense retriever (BGE), and a large language model-based retriever (Qwen3-Embedding). As illustrated in Figure 7, our MemWeaver framework consistently outperforms the standard RAG baseline across all tested embedding models. On the LaMP-1 classification task, MemWeaver achieves a significant improvement in accuracy, and on the LaMP-5 generation task, it consistently yields higher ROUGE-1 scores. This demonstrates that the architectural advantages of our method—namely, its ability to structurally model user history—provide a robust performance lift, independent of the underlying retriever’s strength. The gains are therefore attributable to the novel memory graph structure itself, rather than being an artifact of a single, powerful embedding model.
D.2. Impact of Base Model Scaling
@@
To analyze the impact of the base model’s scale on our framework’s performance, we conducted an experiment to verify that the improvements over the RAG baseline are consistent across models of varying sizes. We evaluated both MemWeaver and the RAG baseline using the Qwen3-3B, Qwen3-8B, Qwen3-14B, and Qwen3-32B models. The results, presented in Figure 8, clearly show two trends. First, both methods benefit from larger base models, with performance on both LaMP-1 (Accuracy) and LaMP-5 (ROUGE-1) improving as the model size increases. Second, and more importantly, our MemWeaver framework maintains a stable and significant performance advantage over the RAG baseline across all tested model sizes. This demonstrates that the benefits of our structured memory approach are complementary to the inherent capabilities of the underlying LLM. The consistent performance gap suggests that MemWeaver’s architectural improvements provide a foundational enhancement that scales effectively with the base model’s power.

### v2

@@ 4.4.2. Analysis of Incremental Update Strategy
To rigorously evaluate the practicality and scalability of our framework in dynamic environments, we designed a simulation experiment modeling the sequential arrival of user interaction data. As illustrated in Figure 5, we benchmark our proposed incremental update strategy against two boundary conditions: a Full Rebuild approach and a No Update baseline. The Full Rebuild method establishes the theoretical accuracy ceiling by utilizing the complete data context for every update; however, its prohibitive computational cost renders it unsustainable for real-time applications. Conversely, the No Update strategy represents the efficiency ceiling but suffers from severe performance degradation, as it fails to capture evolving user interests with a stale profile. Our incremental strategy successfully navigates this trade-off, offering a robust solution that combines the strengths of both extremes. In terms of accuracy, it closely tracks the optimal performance of the Full Rebuild, with any minor initial gaps diminishing rapidly as the user history matures and the graph becomes more complete. Crucially, this high fidelity is achieved with a computational overhead that remains minimal, comparable to the near-instantaneous No Update baseline. These results confirm that our mechanism enables continuous user profile refinement without the latency bottlenecks of graph reconstruction, making it highly suitable for streaming personalization systems.
@@ 5. Conclusion
In this study, we propose MemWeaver, a novel framework that addresses the challenge of deep personalization from unstructured user histories by restructuring them into a hierarchical memory. The framework’s core innovation is the effective interaction between its two memory components. The behavioral memory grounds the model in specific, contextual actions, and the cognitive memory provides abstract guidance on long-term preferences. Crucially, the framework also supports efficient incremental updates, ensuring its scalability and responsiveness in dynamic applications. Extensive experiments on the six datasets of the LaMP benchmark demonstrate the efficacy of this approach. We hope our framework offers new inspiration for the future of user behavior modeling.
Acknowledgments
This research was supported by grants from the National Natural Science Foundation of China (U25B2072, 62502486), the Key Technologies R & D Program of Anhui Province (No. 202423k09020039) and the Fundamental Research Funds for the Central Universities.
@@ Appendix B Dataset Details
Table 4. Performance on Diverse Benchmarks.
Dataset Metric Non-P MemoRAG Memory-R1 CFRAG MemWeaver
OpinionQA Acc. 0.352 0.397 0.423 0.452 0.473
AmazonRev R-1 0.221 0.263 0.246 0.286 0.302
ChangeMyV Avg. 26.59 26.98 30.15 31.25 35.33
Table 5. Comparison with Memory-based Baselines on LaMP.
Method LaMP-1 LaMP-2 LaMP-3 LaMP-4 LaMP-5 LaMP-7
(Acc) (Acc) (MAE) (R-1) (R-1) (R-1)
Memory-R1 0.579 0.385 0.398 0.156 0.454 0.440
MemoRAG 0.650 0.392 0.344 0.163 0.470 0.452
Mem0 0.652 0.424 0.330 0.162 0.465 0.451
MemWeaver 0.673 0.463 0.280 0.172 0.475 0.479
@@
Appendix C Additional Experimental Results
In this section, we provide extended experimental results to further validate the robustness and superiority of the MemWeaver framework. This includes evaluations on diverse benchmarks, comparisons with state-of-the-art memory-augmented LLMs, and sensitivity analyses regarding embedding models and backbone scaling.
C.1. Robustness Across Diverse Benchmarks
To verify the generalizability of MemWeaver beyond the LaMP benchmark, we extend our evaluation to three additional datasets characterized by distinct personalization challenges: OpinionQA (response prediction) (Sun et al., 2025), Amazon Reviews (user-specific sentiment) (Qiu et al., 2025), and ChangeMyView (persuasive reasoning) (Zhang et al., 2025c). As shown in Table 4, MemWeaver consistently surpasses competitive baselines (including CFRAG and Memory-R1) across all metrics. These results demonstrate that our hierarchical memory structure effectively captures user-specific nuances across various domains, from subjective opinions to stylistic preferences.
C.2. Comparison with Memory-based Baselines
We further compare MemWeaver with recent memory-augmented LLM frameworks, including MemoRAG (Qian et al., 2025), Memory-R1 (Yan et al., 2025), and Mem0. While these baselines focus on long-context compression or simple key-value storage, they often neglect the structured correlations within user history. As summarized in Table 5, MemWeaver outperforms these specialized memory models across all LaMP tasks. The performance gap highlights the necessity of our dual-component design, which jointly models behavioral contexts and cognitive abstractions.
C.3. Impact of Different Embedding Models
@@
To verify that the performance gains of our framework are not solely dependent on a specific embedding model, we evaluate its robustness across three distinct retrieval backbones: a sparse retriever (BM25), a strong dense retriever (BGE), and a large language model-based retriever (Qwen3-Embedding). As illustrated in Figure 7, MemWeaver consistently outperforms the standard RAG baseline regardless of the retriever used. This demonstrates that the architectural advantages of our method—namely, its ability to structurally model user history—provide a robust performance lift independent of the underlying retriever’s strength.
C.4. Impact of Base Model Scaling
@@
Finally, we analyze the impact of the base model’s scale by evaluating MemWeaver using Qwen3-3B, 8B, 14B, and 32B. As presented in Figure 8, performance on both LaMP-1 and LaMP-5 improves as model size increases. More importantly, MemWeaver maintains a stable and significant advantage over the RAG baseline across all tested scales. This consistency suggests that our structured memory approach provides a foundational enhancement that scales effectively with the base model’s inherent capabilities.
Appendix D Prompt Templates
This section details the prompt structures in MemWeaver. The design prioritizes consistency and interpretability by explicitly structuring roles, contexts, and instructions to standardize how the LLM extracts and applies user-level knowledge.
D.1. Prompt for Cognitive Memory
The cognitive memory is constructed via a two-stage summarization pipeline that abstracts long-term user preferences from fine-grained history to coarse-grained semantics.
In the Local Summary stage, the model acts as an ”expert at analyzing user behavior patterns” to summarize activity clusters (formatted as Cluster X Activities (Y records):). It extracts key themes and local coherence while filtering noise. Subsequently, in the Global Summary stage, the model functions as an ”expert at creating concise user preference summaries.” It aggregates all local summaries (Cluster N: [Text]) into a unified profile (max 300 words). This hierarchical approach ensures semantic generalization by synthesizing specific behaviors into enduring cognitive traits.
D.2. Prompt for Personalized Generation
MemWeaver employs a unified personalized generation prompt that integrates two memory sources: cognitive memory (high-level global priors) and behavioral memory (local, example-level cues).
As shown in Table 6, this modular structure places both memory components before the task-specific query. By providing complementary signals—global abstractions and concrete evidence—the prompt steers the model toward user-aligned reasoning. This design enhances controllability and transferability across diverse personalized tasks within the LaMP benchmark.